OpenStudy (anonymous):

find the derivative of f(x)=(2x+1)^3(3x-2)^2 and simplify to a single term. 4(2x+1)^3 (3x-2) + 9(2x+1)^2 (3x-2)^2 I have this so fare.

OpenStudy (anonymous):

combine the like terms ( 2x+1) and (3x-2) ans simplify

OpenStudy (anonymous):

if you multiply it all out you will get a function involving powers of x and a constant. But notice that this is a function of of ( 2x+1) and (3x-2), you can probably factor the binomial and make this a bit easier.

OpenStudy (anonymous):

say you want to simplify: 4(2x+1)^3 (3x-2) By the binomial theorem, (2x+1)^3=8x^3+12x^2+6x+1 (8x^3+12x^2+6x+1)(3x-2)=24x^5+20x^3-6x^2-9x-2 that's the model for the first, now do the second one. it takes a long time, but you just have to be patient and meticulous. :)

OpenStudy (zarkon):

Factor out \[(2x+1)^2\] and \[3x-2\] you will end up with \[6(2x+1)^2(3x-2)(5x-1)\]

OpenStudy (anonymous):

easier method use u and v u=2x+1 v=3(3x-2)^2

OpenStudy (anonymous):

so then its simply u^v

OpenStudy (anonymous):

In the formula you posted you have the coefficients wrong. It looks like you got confused about what derivative to multiply by when doing the chain rule. You should get 6 for both coefficients. Zarkon's answer is correct.